Managing activities across the value chain represents a comprehensive integrated approach to the traditional management functions of planning and control. Eliminating non-value-added activities from the chain is central to this strategic approach to cost management.

Smith Corporation is considering the implementation of a JIT inventory system. The company recently analyzed its cycle time to determine the average number of days spent in each activity of its production process.
Answer #1

Part a)

The steps that are needed to determine the value-adding production activities are given as follow:

1) Check that the activity contributes in the conversion of an item/raw material into a usable end product. (such as cutting materials, assembling materials, etc.)

2) Check that the customer is willing to pay for the activity. (a customer would be interested in the final product and would prefer to pay only for activities that are essential for getting the raw material converted into finished good)

3) Check that the proposed activity is performed in the correct way the very first time it is carried out.

_____

Part b)

The total cycle time in days is the sum of number of days taken by each activity and is determined as below:

Production Activity Number of Days
Receiving Materials 1
Inspecting Materials 4
Storing Materials 9
Moving Materials into Production 6
Setting-up Production Equipment 7
Cutting Materials 6
Assembling Materials 7
Painting Finished Products 4
Packaging Finished Products 1
Total Cycle Time 45 days

The total cycle time is 45 days.

_____

Part c)

The manufacturing efficiency ratio is calculated as follows:

Manufacturing Efficiency Ratio = Sum of Days taken by Value-Adding Activities/Total Cycle Time*100

Here, Sum of Days Taken by Value Adding Activities = Cutting Materials + Assembling Materials + Painting Finished Products + Packaging Finished Products = 6 + 7 + 4 + 1 = 18 days and Total Cycle Time = 45 days

Substituting these values in the above formula, we get,

Manufacturing Efficiency Ratio = 18/45*100 = 40%

_____

Part d)

The activities that might be reduced or eliminated to implement an efficient JIT system are listed below:

1) Receiving Materials

2) Storing Materials

3) Moving Materials into Production